# An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data

### v2

@@ Abstract
The estimation of underwater sound velocity distribution serves as a critical basis for facilitating effective underwater communication and precise positioning, given that variations in sound velocity influence the path of signal transmission. Conventional techniques for the direct measurement of sound velocity, as well as methods that involve the inversion of sound velocity utilizing acoustic field data, necessitate on–site data collection. This requirement not only places high demands on device deployment, but also presents challenges in achieving real-time estimation of sound velocity distribution. In order to construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, we propose a self-attention embedded multimodal data fusion convolutional neural network (SA-MDF-CNN) for real-time underwater sound speed profile (SSP) estimation. The proposed model seeks to elucidate the inherent relationship between remote sensing sea surface temperature (SST) data, the primary component characteristics of historical SSPs, and their spatial coordinates. This is achieved by employing CNNs and attention mechanisms to extract local and global correlations from the input data, respectively. The ultimate objective is to facilitate a rapid and precise estimation of sound velocity distribution within a specified task area. The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability, exhibiting reduced error rates and enhanced resistance to disturbances when benchmarked against existing advanced techniques.
@@ I Introduction
Underwater sound velocity construction plays an indispensable role in many applications such as underwater positioning, navigation, timing, communication (PNTC) and target recognition due to its decisive factor in the signal propagation mode [1, 2, 3]. Acoustic propagation speed in marine settings exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7]. Generally, the sound velocity exhibits a stratified variation with depth, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
@@
The methodology for acquiring SSPs has been a popular research topic for a long period. In traditional ways, SSPs can be directly measured using various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP). Alternatively, SSPs can be derived by inversion based on acoustic field data, which encompasses parameters such as signal transmission time or received signal strength [8, 9, 10, 11, 12]. In general, instrument-based measurements can provide precise SSPs with high–depth resolution, but the measurement process is notably time-intensive. For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters to complete the deployment and retrieval of the equipment [11]. As a result, it is difficult to provide real-time sound velocity distributions for underwater communication and positioning through instrument-based measurement. Considering that the distribution of sound velocity can affect the distribution of sound field, researchers have proposed a series of methods to invert the sound velocity distribution by using on-site measured sound field data such as signal propagation time in order to accelerate the process of obtaining sound velocity. There are three main frameworks for inversion methods, including matching field processing (MFP) [13], compressive sensing (CS) [14, 15], and deep learning (DL) [4, 16]. The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional conductivity-temperature-depth (CTD) sensors and sound velocity profilers (SVPs) in marine data collection. However, because of the interference of noise in sound field measurement, there is a loss in the accuracy of these sound velocity inversion methods. More importantly, it still takes a certain amount of time to measure the sound field data, and new requirements have been put forward for the deployment of expensive sound field measurement equipment. The implementation of SSP inversion techniques is inherently constrained by their dependence on active sonar measurements, thereby limiting their applicability in regions beyond the operational range of underwater acoustic monitoring networks.
@@
Over the past few years, the fast development of underwater automated sensing equipment has provided increasingly rich vertical observation reference data for exploring the distribution of underwater sound velocity. To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use empirical sound velocity distribution data or remote sensing data. [17, 18, 19] utilize empirical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas. These SSP prediction methods have a single type of input data and low complexity of model construction. However, because of the low time resolution of historical samples, its accuracy performance is not good enough in estimating the sound velocity within a small-scale time range. In fact, the variation of sound velocity distribution in a small-scale spatial area is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes. Marine remote sensing technology provides real-time and reliable high spatiotemporal resolution sea surface temperature (SST) data [20], which can offer initial sea surface condition constraints for estimating underwater sound velocity distribution. [21] proposed a self-organizing map (SOM) neural network model that combines empirical sound velocity data and remote sensing SST data, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution. As a result, the accuracy performance has not been significantly improved compared to prediction methods.
@@
In order to fast and accurately estimate the SSP of a given task area without on-site underwater data measurement, we propose an interpretable self-attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model in this paper to deeply capture the intrinsic relationships among empirical SSP data, remote sensing SST data, and spatial coordinates. The core idea is to capture the local feature correlation of regional sound velocity distribution through CNN, and enhance the learning ability of global feature correlation of multimodal data through attention mechanism. To evaluate the effectiveness of the proposed SA-MDF-CNN model, we conducted experiments using Argo data from the Pacific Ocean and measured data from the South China Sea in 2023. Results show that SA-MDF-CNN can achieve lower root mean square error (RMSE) than other spatial SSP construction methods. The contribution of this paper can be summarized as follows:
@@
  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model. We take into account the mutual influence among empirical sound velocity distribution, SST, and spatial position, and the attention module is designed to enhance the global feature learning ability of the model.
@@
  • To enhance the interpretability of the model, the weights of the model parameters are visualized. As the number of training iterations increases, the model parameter weights gradually focus on the shallow water part. This is reasonable because through the empirical orthogonal function (EOF) decomposition of historical SSPs, it can be seen that the differences in sound velocity distribution are more prominent in shallow water. Therefore, the weight visualization results indicate that the model can effectively capture the characteristics of sound velocity distribution.
The structure of this article is arranged as follows. Section 2 provides a detailed introduction to relevant works. The overall architecture and functional module description of the SA-MDF-CNN model are provided in Section 3. Section 4 gives the experimental findings, and the final section, Section 5, presents the concluding remarks.
@@ II Related works
The distribution of underwater sound velocity plays an essential role in the energy distribution and propagation trajectory of acoustic signals. Therefore, real-time and accurate estimation of sound velocity distribution has significant value for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
@@
The traditional way of obtaining sound velocity distribution is usually through direct measurement using shipborne CTD or SVP equipment, which has the advantage of high accuracy [22, 23]. However, it comes with high economic costs and measurement time expenses. With the development of sensing and network technology, many kinds of underwater environmental observation systems have been established to enhance human understanding of the ocean. In 1979, MUNK and Wunsch first put forward the concept of ocean acoustic tomography that using acoustic field observation information to invert the sound velocity distribution [24, 25]. Since then, there have been three mainstream frameworks for sound velocity inversion, namely MFP [13], CS [14, 15], and DL [4, 16]. The MFP framework mainly consists of four steps. Firstly, EOF decomposition is adopted to extract the principal component features of the regional sound velocity distribution. Then, different feature combinations are generated to form candidate SSPs. Next, the sound field distribution is simulated based on ray theory. Finally, the simulated sound field is matched with the measured sound field to determine the optimal estimation of SSP. To accelerate the search speed of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but resulted in suboptimal solutions [13]. Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion, but they also obtained suboptimal solutions [26, 27]. To further improve the efficiency of inversion algorithm execution, Bianco [15] and Choo [14] proposed SSP inversion framework based on CS, respectively. In the CS framework, the mapping relationship from the sound field distribution to the sound velocity distribution is established through a matrix, which eliminates the search process for matching terms. However, the matrix relationship introduces linear approximation, thus sacrificing inversion accuracy.
@@
In recent years, deep learning theory has developed rapidly and achieved significant results in constructing unknown nonlinear relationships between different data [28, 29]. In addition, long-term underwater environmental observations have accumulated a large amount of data for marine hydrological research, laying a data foundation for the application of deep learning underwater. To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, we proposed a auto-encoding feature mapping neural network model for SSP inversion in our early work [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion. Considering the insufficient accumulation of historical sound velocity data in some ocean areas, the DL model is prone to over-fitting and reduces accuracy performance. Therefore, we further propose a few-shot meta learning framework to accelerate the convergence speed of the model [16]. However, the above-mentioned methods based on MFP, CS, and DL all rely on real-time measured acoustic field data, which puts high demands on the deployment of underwater observation equipment. Therefore, these methods not only face high equipment economic costs, but also have limited application scope for areas without sonar measurement systems.
@@
Nowadays, how to eliminate the need for underwater on-site data measurement has become a research hotspot in the field of SSP inversion. Based on this requirement, Liu et al. and Lu et al. both proposed a long-short term memory (LSTM) neural network model for SSP prediction [17, 18], which only requires empirical SSP data. However, due to the low temporal resolution of prior data in most marine areas, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results. In fact, the SSP estimation methods, that rely on a single data modality (sound field data or historical SSP data), are susceptible to data quality issues, such as high sound field measurement noise or low sound velocity sampling time resolution. To improve the robustness of the sound velocity estimation model, researchers have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21]. Yu et al. proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data. Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, making it difficult to accurately describe the spatiotemporal distribution changes of sound velocity. Xu et al. proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data. However, SOM can only capture the influence of sea surface temperature on surface sound velocity distribution, and lacks the ability to capture large-scale feature correlations. Therefore, the accuracy of sound velocity estimation is difficult to meet practical application requirements.
@@ III SA-MDF-CNN Structure for SSP Estimation
To realize real-time and accurate estimation of sound velocity distribution, we propose an SSP estimation structure based on SA-MDF-CNN, which is shown in figure 1. In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research scope. The remote sensing SST data, latitude and longitude coordinates, and the first three order feature vectors of historical SSPs decomposed by EOFs are first fused to construct the fusion data as the SA-MDF-CNN training data. Then, the multimodal fusion data of 8 grids around the task region is combined to construct the target SSP through the proposed SA-MDF-CNN model. In the following part, we will provide detailed introductions to data sources, data fusion structures, and the composition of neural network model, separately.
@@ III-A Data sources
Remote sensing SST data are provided by the National Oceanic and Atmospheric Administration (NOAA) [31] with a spatial grid resolution of 0.25 and a temporal resolution of 1 day. The SSP data are obtained from the Chinese Observation and Research Station fo Global Ocean Argo System (Hangzhou) [32] with a spatial grid resolution of 1 and a temporal resolution of 1 month. Based on the objective analysis method of gradient-dependent correlation scale optimal interpolation, the 3D grid data of the subsurface layer (5-2000 meters) that covering 179.5W to 179.5E and 89.5S to 89.5N was constructed, and the observation profile was vertically interpolated to 57 standard layers with unequal intervals.

### v3

# An Attention-Assisted Multi-Modal Data Fusion Model for Real-Time Estimation of Underwater Sound Velocity

@@ Abstract
The estimation of underwater sound velocity distribution serves as a critical basis for facilitating effective underwater communication and precise positioning, given that variations in sound velocity influence the path of signal transmission. Conventional techniques for the direct measurement of sound velocity, as well as methods that involve the inversion of sound velocity utilizing acoustic field data, necessitate on–site data collection. This requirement not only places high demands on device deployment, but also presents challenges in achieving real-time estimation of sound velocity distribution. In order to construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, we propose a self-attention embedded multimodal data fusion convolutional neural network (SA-MDF-CNN) for real-time underwater sound speed profile (SSP) estimation. The proposed model seeks to elucidate the inherent relationship between remote sensing sea surface temperature (SST) data, the primary component characteristics of historical SSPs, and their spatial coordinates. This is achieved by employing CNNs and attention mechanisms to extract local and global correlations from the input data, respectively. The ultimate objective is to facilitate a rapid and precise estimation of sound velocity distribution within a specified task area. The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability against existing advanced techniques.
@@ I Introduction
Underwater sound velocity construction plays an indispensable role in many applications such as underwater positioning, navigation, timing, communication (PNTC) and target recognition due to its decisive factor in the signal propagation mode [1, 2, 3]. Acoustic propagation speed in marine environment exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7]. Generally, the sound velocity exhibits a stratified variation with depth within the kilometer level range, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
@@
The methodology for acquiring SSPs has been a popular research topic for a long period. In traditional ways, SSPs can be directly measured by various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP). Alternatively, SSPs can be derived by inversion based on acoustic field data, which encompasses parameters such as signal transmission time or received signal strength [8, 9, 10, 11, 12]. In general, instrument-based measurements can provide precise SSPs with high–depth resolution, but the measurement process is notably time-intensive. For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters, including the deployment and retrieval of the equipment [11]. As a result, it is difficult to provide real-time sound velocity distributions for underwater communication and positioning through instrument-based measurement.
Since the distribution of the sound speed affects the distribution of the acoustic field, scholars in related areas have come up with a range of approaches to invert the sound speed distribution. They make use of the on-site measured data of the sound field such as the signal propagation time, aiming at accelerating the process of obtaining the sound velocity distribution. There are three main frameworks for inversion methods, including matching field processing (MFP) [13], compressive sensing (CS) [14, 15], and deep learning (DL) [4, 16]. The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional CTD sensors and SVPs in marine data collection. However, because of the interference of noise in sound field measurement, there is a loss in the accuracy of these sound velocity inversion methods. More importantly, it still takes a certain amount of time to measure the sound field data, and new requirements have been put forward for the deployment of expensive sound field measurement equipment. The implementation of SSP inversion techniques is inherently constrained by their dependence on active sonar measurements, thereby limiting their applicability in regions beyond the operational range of underwater acoustic monitoring networks.
@@
Over the past few years, the fast development of underwater automated sensing equipment has provided increasingly rich vertical observation reference data for exploring the distribution of underwater sound velocity. To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use historical sound velocity distribution data or remote sensing data. [17, 18, 19] utilize historical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas. These SSP prediction methods have a single type of input data and low complexity of model construction. However, because of the low time resolution of historical samples, its accuracy performance is not good enough in estimating the sound velocity within a small-scale time range. In fact, the temporal variation of sound velocity distribution in small-scale spatial regions is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes. Marine remote sensing technology provides real-time and reliable high spatiotemporal resolution sea surface temperature (SST) data [20], which can offer initial sea surface condition constraints for estimating underwater sound velocity distribution. Xu et al [21] introduced an innovative self-organizing map (SOM) architecture that integrates empirical acoustic propagation measurements with satellite-derived SST observations to establish robust correlations between marine acoustic properties and surface thermal patterns, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution. As a result, the accuracy performance is difficult to meet the application requirements of high-precision PNTC systems.
@@
In order to fast and accurately estimate the SSP of a given task area without on-site underwater data measurement, we propose an interpretable self-attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model in this paper to deeply capture the intrinsic relationships among empirical SSP data, remote sensing SST data, and spatial coordinates. The core idea focuses on extracting spatial patterns of regional sound velocity interdependencies via convolutional neural networks, while the application of attention mechanisms enhances the model’s capacity to learn complex interactions within multimodal data. In an attempt to comprehensively appraise the overall effectiveness of the purportedly proposed SA-MDF-CNN model, we carried out a series of carefully designed experiments by utilizing the Argo data originating from the expanse of the Pacific Ocean and the precisely measured data obtained from the South China Sea in the year 2023. The detailed results clearly indicate that the SA-MDF-CNN model has the distinct ability to accomplish a notably lower root mean square error (RMSE) value compared to various other existing spatial SSP construction methods. The significant contributions presented in this research paper can be summarized in the following ways:
@@
  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an interpretable SA-MDF-CNN model. We take into account the mutual influence among historical sound velocity distribution, SST, and spatial position, and the attention module is developed to optimize the model’s comprehensive feature extraction capabilities across spatial domains.
@@
  • To boost the model’s interpretability, a visualization of the model parameter weights is carried out. With the escalation of training iterations, the weights of model parameters progressively concentrate on the shallow water region. This phenomenon is justifiable as, upon performing empirical orthogonal function (EOF) decomposition on historical SSPs, it becomes evident that disparities in sound velocity distribution are more pronounced in shallow water. Consequently, the outcomes of weight visualization suggest that the model is capable of efficiently capturing the traits of sound velocity distribution.
The organization of this paper is designed as follows. In Section 2, a brief elaboration on related works is offered. Section 3 puts forward the general framework and the descriptive details of the functional modules of the SA-MDF-CNN. The outcomes of the experiment are reported in Section 4. Concluding statements are presented in Section 5.
@@ II Related works
The spatial structure of underwater sound velocity field plays a decisive role in the energy distribution and propagation trajectory of acoustic signals. Therefore, real-time and accurate estimation of sound velocity distribution is of great meaningful for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
@@
The traditional way of obtaining sound velocity distribution is usually through direct measurement using shipborne CTD or SVP equipment, which has the advantage of high accuracy [22, 23]. However, it comes with high economic costs and measurement time expenses. The development of sensor network technologies has led to the implementation of multiple underwater environmental monitoring frameworks, advancing scientific comprehension of oceanic processes. The methodological framework of underwater acoustic inversion was initially formalized in 1979 by Munk et al.[24, 25], introducing an innovative paradigm that leverages hydroacoustic propagation measurements for reconstructing SSPs. Since then, there have been three mainstream frameworks for sound velocity inversion, namely MFP [13], CS [14, 15], and DL [4, 16]. The MFP framework mainly consists of four steps. Firstly, EOF decomposition is adopted to extract the principal component features of the regional sound velocity distribution. Then, different feature combinations are generated to form candidate SSPs. Subsequently, acoustic field distributions are computationally modeled employing an acoustic ray tracing theory. The concluding phase involves cross-validation between the modeled sound field distributions and experimentally measurements to establish the optimal inversion parameters for the SSP. To accelerate the search process of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but only suboptimal solutions were obtained [13]. Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion [26, 27], but they have the same problem as [13]. To further improve the efficiency of inversion algorithm execution, Bianco [15] and Choo [14] proposed SSP inversion framework based on CS, respectively. Within the context of the CS framework, the correlation between the sound field distribution and the sound velocity distribution is formulated via a matrix based approach, which eliminates the search process for matching terms. However, the matrix relationship introduces linear approximation, thus sacrificing inversion accuracy.
@@
The theoretical frameworks of deep neural networks (DNNs) have experienced accelerated progression in contemporary research, particularly in establishing intricate nonlinear dependencies across heterogeneous data domains [28, 29]. In addition, long-term underwater environmental observations have accumulated a large amount of data for marine hydrological research, laying a data foundation for the application of deep learning underwater. To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, an auto-encoding neural network model for feature mapping was introduced for SSP inversion in our prior research [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion. Considering the insufficient accumulation of historical sound velocity data in some ocean areas, the DL model is prone to over-fitting and reduces accuracy performance. Therefore, we proposed a meta-learning framework tailored for few-shot scenarios to enhance the rate at which the model converges [16]. However,the methodologies aforementioned, which are predicated on MFP, CS, and DL, all rely on real-time measured acoustic field data, which imposes strict requirements on the arrangement of underwater observation equipment. Therefore, these methods not only face high equipment economic costs, but also have limited application scope for areas without sonar measurement systems.
@@
Nowadays, removing the necessity for on-site underwater data measurement has emerged as a prominent research focus in the domain of SSP inversion. In accordance with this demand, both Liu et al. [17] and Lu et al. [18] put forward a Long-Short Term Memory (LSTM) neural network model for SSP forecasting, which solely requires historical SSP data. Nevertheless, owing to the insufficient temporal resolution of prior data in the majority of marine regions, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results. In fact, the SSP estimation methods, that rely on a single data modality (sound field data or historical SSP data), are susceptible to data quality issues, such as high sound field measurement noise or low sound velocity sampling time resolution. To improve the robustness of the sound velocity estimation model, scholars have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21]. Yu et al. [30] proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data. Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, resulting in challenges in accurately depicting the spatiotemporal variations of sound velocity. Xu et al. [21] proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data. However, SOM is confined to capturing the impact of SST on the distribution of sound velocity at the surface, and lacks the ability to capture large-scale feature correlations. Therefore, the accuracy of sound velocity estimation is difficult to meet practical application requirements.
@@ III SA-MDF-CNN Structure for SSP Estimation
To realize real-time and accurate estimation of sound velocity distribution, we propose an SSP estimation structure based on SA-MDF-CNN, which is shown in figure 1. In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research area. The remote sensing SST data, latitude and longitude coordinates, and principal component eigenvectors derived from EOF decomposition of historical SSPs are first fused to construct the fusion data as the SA-MDF-CNN training data. Then, the multimodal fusion data of 8 grids around the task region are combined through the proposed SA-MDF-CNN model to construct the target SSP. In the following part, we will provide detailed introductions to data sources, data fusion structures, and the composition of neural network model, separately.
@@ III-A Data sources
Remote sensing SST data are provided by the National Oceanic and Atmospheric Administration (NOAA) [31] with a spatial grid resolution of 0.25 and a temporal resolution of 1 day. The SST data are arithmetic averaged on a monthly basis to achieve the consistency of the fused data in time sampling. The SSP data are obtained from the Chinese Observation and Research Station fo Global Ocean Argo System (Hangzhou) [32] with a spatial grid resolution of 1 and a temporal resolution of 1 month. Based on the objective analysis method of gradient-dependent correlation scale optimal interpolation, the 3D grid data of the subsurface layer (5-2000 meters) that covering 179.5W to 179.5E and 89.5S to 89.5N was constructed, and the observation profile was vertically interpolated to 57 standard layers with unequal intervals.